A charge of 6.04 mC is placed at each corner of a square 0.180 m on a side. Determine the magnitude of the force on each charge.
Answer #1
Concepts and reason

The concept of electric force is required to solve the problem.

First, determine the length of the diagonal of the square. Then determine the force on one of the charge due to each remaining charge. Then determine the direction of each force and find out the magnitude of net force on that charge.

Fundamentals

Electrostatic force is the force between two charges separated by a distance r. The force on charge q2{q_2} due to the charge q1{q_1} is given as,

F=kq1q2r122F = \frac{{k{q_1}{q_2}}}{{{r_{12}}^2}}

Here, k is the coulomb’s constant and r12{r_{12}} is the distance between two charges.

The figure given below shows the arrangement of charges on the corners of square of side a.

719
a
F24
F4

The force on charge 4 due to charge 1 is,

F14=kq1q4r142{F_{14}} = \frac{{k{q_1}{q_4}}}{{{r_{14}}^2}}

Substitute q for q1{q_1} and q4{q_4} , and a for r14{r_{14}} in the above equation.

F14=kq2a2{F_{14}} = \frac{{k{q^2}}}{{{a^2}}}

The force on charge 4 due to charge 3 is,

F34=kq3q4r142{F_{34}} = \frac{{k{q_3}{q_4}}}{{{r_{14}}^2}}

Substitute q for q3{q_3} and q4{q_4} , and a for r14{r_{14}} in the above equation.

F34=kq2a2{F_{34}} = \frac{{k{q^2}}}{{{a^2}}}

Determine the distance between the charge 2 and charge 4 using Pythagoras theorem.

r24=a2+a2=2a\begin{array}{c}\\{r_{24}} = \sqrt {{a^2} + {a^2}} \\\\ = \sqrt 2 a\\\end{array}

The force on charge 4 due to charge 2 is,

F24=kq2q4r242{F_{24}} = \frac{{k{q_2}{q_4}}}{{{r_{24}}^2}}

Substitute 2a\sqrt 2 a for r24{r_{24}} and q for q2{q_2} and q4{q_4} in the above equation.

F24=kq22a2{F_{24}} = \frac{{k{q^2}}}{{2{a^2}}}

Resolve the force F24{F_{24}} into horizontal and vertical component.

F24x F34 4 so
19
a
F24
F247

The horizontal component of force F24{F_{24}} is,

F24x=F24cosθ{F_{24x}} = {F_{24}}\cos \theta

The vertical component of force F24{F_{24}} is,

F24y=F24sinθ{F_{24y}} = {F_{24}}\sin \theta

The angle θ\theta is given as,

θ=tan1(aa)=45o\begin{array}{c}\\\theta = {\tan ^{ - 1}}\left( {\frac{a}{a}} \right)\\\\ = {45^{\rm{o}}}\\\end{array}

Refer the above figure and determine the net force in the horizontal direction.

Fx=F34+F24cosθ{F_x} = {F_{34}} + {F_{24}}\cos \theta

Substitute kq2a2\frac{{k{q^2}}}{{{a^2}}} for F34{F_{34}} , kq22a2\frac{{k{q^2}}}{{2{a^2}}} for F24{F_{24}} , and 45o{45^{\rm{o}}} for θ\theta in the above equation.

Fx=kq2a2+kq22a2cos45o=kq2a2(1+122)\begin{array}{c}\\{F_x} = \frac{{k{q^2}}}{{{a^2}}} + \frac{{k{q^2}}}{{2{a^2}}}\cos {45^{\rm{o}}}\\\\ = \frac{{k{q^2}}}{{{a^2}}}\left( {1 + \frac{1}{{2\sqrt 2 }}} \right)\\\end{array}

Determine the net force in the vertical direction.

Fy=F14+F24sinθ{F_y} = {F_{14}} + {F_{24}}\sin \theta

Substitute kq2a2\frac{{k{q^2}}}{{{a^2}}} for F14{F_{14}} , kq22a2\frac{{k{q^2}}}{{2{a^2}}} for F24{F_{24}} , and 45o{45^{\rm{o}}} for θ\theta in the above equation.

Fy=kq2a2+kq22a2sin45o=kq2a2(1+122)\begin{array}{c}\\{F_y} = \frac{{k{q^2}}}{{{a^2}}} + \frac{{k{q^2}}}{{2{a^2}}}\sin {45^{\rm{o}}}\\\\ = \frac{{k{q^2}}}{{{a^2}}}\left( {1 + \frac{1}{{2\sqrt 2 }}} \right)\\\end{array}

The magnitude of the net force on charge 4 is given as,

Fnet=Fx2+Fy2{F_{{\rm{net}}}} = \sqrt {{F_x}^2 + {F_y}^2}

Substitute kq2a2(1+122)\frac{{k{q^2}}}{{{a^2}}}\left( {1 + \frac{1}{{2\sqrt 2 }}} \right) for Fx{F_x} and Fy{F_y} in the above equation.

Fnet=(kq2a2(1+122))2+(kq2a2(1+122))2=kq2a2(1+122)2\begin{array}{c}\\{F_{{\rm{net}}}} = \sqrt {{{\left( {\frac{{k{q^2}}}{{{a^2}}}\left( {1 + \frac{1}{{2\sqrt 2 }}} \right)} \right)}^2} + {{\left( {\frac{{k{q^2}}}{{{a^2}}}\left( {1 + \frac{1}{{2\sqrt 2 }}} \right)} \right)}^2}} \\\\ = \frac{{k{q^2}}}{{{a^2}}}\left( {1 + \frac{1}{{2\sqrt 2 }}} \right)\sqrt 2 \\\end{array}

Substitute 8.99×109Nm2/C28.99 \times {10^9}{\rm{ N}} \cdot {{\rm{m}}^2}{\rm{/}}{{\rm{C}}^{\rm{2}}} for k, 0.180 m for a, and 6.04 mC for q in the above equation.

Fnet=(8.99×109Nm2/C2)(6.04mC(103C1mC))2(0.180m)2(1+122)2=1.94×107N\begin{array}{c}\\{F_{{\rm{net}}}} = \frac{{\left( {8.99 \times {{10}^9}{\rm{ N}} \cdot {{\rm{m}}^2}{\rm{/}}{{\rm{C}}^{\rm{2}}}} \right){{\left( {6.04{\rm{ mC}}\left( {\frac{{{{10}^{ - 3}}{\rm{ C}}}}{{1{\rm{ mC}}}}} \right)} \right)}^2}}}{{{{\left( {0.180{\rm{ m}}} \right)}^2}}}\left( {1 + \frac{1}{{2\sqrt 2 }}} \right)\sqrt 2 \\\\ = 1.94 \times {10^7}{\rm{ N}}\\\end{array}

As all the charges are equal, the net force on each charge will be same.

Ans:

The magnitude of net force on each charge is 1.94×107N1.94 \times {10^7}{\rm{ N}} .